OpenStudy (anonymous):

Tom has 3 red sweets 3 green sweets 5 yellow sweets What is the portability of him NOT getting green? What is the portability of him NOT getting purple? Thanks

OpenStudy (anonymous):

What is the probability of him getting green?\[ \frac{ 3 }{ 3 + 3 + 5 } = \frac{ 3 }{ 11 }\] The probability of him NOT getting green is \[1 - \frac{ 3 }{ 11 } = \frac{ 8 }{ 11 }\] OR The probability of him NOT getting green is nothing but the probability of him getting (Red + Yellow) Therefore, it would be \[\frac{ 3 + 5 }{ 3 + 5 + 3 } = \frac{ 8 }{ 11 }\] Now, what is the probability of him getting purple? Zero. Why? Because there are no purple sweets. \[\frac{ 0 }{ 11 } = 0\]

OpenStudy (anonymous):

Thanks thats helped a lot thanks. So for sweets where there are none its 0/(AMOUNT OF SWEETS) In this case 0/11 thanks